A cylinder has a diameter of 12 in. and a height of 6 in. Find the surface area.
Answer by TimothyLamb(4379)

a = 2pi*r*h + 2pi*rr
r = d/2 = 12/2 = 6 in
h = 6 in
---
a = 2*3.1415*6*6 + 2*3.1415*6*6
a = 452.376 sq.in
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php